Find the emf induced in a 0.1-H coil when the current in it is changing at the rate of 80 A/s.


1
Expert's answer
2021-11-04T18:07:14-0400

The expression for the induced emf:

"E = -L \times \frac{di}{dt}"

Inductance of the coil= L = 0.1 H

Rate of change of the current in the coil "= \frac{di}{dt} = 80 \;A/s"

"E = -0.1 \times 80 \; Volt = -8.0 \; V"

Therefore, induced emf in the coil = 8.0 Volt